OpenStudy (anonymous):

pvc cable 32mm sqare space factor 40by60 and truncking50by25the number of cable

OpenStudy (triciaal):

welcome to openstudy!! would love to help but have no idea what you mean by what is posted. please write the question in English.

OpenStudy (anonymous):

Is it a 32 mm diameter cable?

OpenStudy (mathmate):

These are calculations that relate to local codes, so answers depend on where the cables are to be installed. Also, trunking factor depend on the length of the run, as well as the number of bends. Following calculations are approximate, and provided to give a general idea only for a short straight run. PVC cable 32 \(mm^2\) has an approximate cable factor of 90. Trunking factor for 40x60 is about 1000, so number of cables = 1000/90=11 approx. Trunking factor for 50x25 is about 500, so number of cables = 500/90=5 approx. Note: the cables factors and trunking factors are approximate. For more information, you can see: http://www.rm-electrical.com/publish/technical-specs/files/2011/10/Conduit-Trunking-Capacity-Chart_RM-Technical.pdf http://www.tlc-direct.co.uk/Book/4.5.3.htm https://books.google.ca/books?id=Kj1HInMT0LkC&pg=PA86&lpg=PA86&dq=cable+factor+32+mm2&source=bl&ots=FIFzqWbomN&sig=Xtn9fF8mVfsEphegBwMHBEFGYW4&hl=en&sa=X&ei=G56lVLSQOoS3yATK3YGoAw&ved=0CB8Q6AEwAA#v=onepage&q=cable%20factor%2032%20mm2&f=false